# A Survey on Deep Learning based Time Series Analysis with Frequency Transformation

### v5

@@ 3.2. Compression
Previous works utilize frequency transformation to obtain sparse representations and remove redundant information in the frequency domain. Moreover, since noise signals usually appear as high frequencies, it is easy to filter them out in the frequency domain. For example, in FiLM (FiLM_2022, ), authors view time series forecasting from the sequence compression perspective and apply Fourier analysis to keep the part of the representation related to low-frequency Fourier components to remove the impact of noises. (Rippel_nips_2015, ) proposes spectral pooling that performs dimensionality reduction by truncating the representation in the frequency domain because energy is heavily concentrated in the lower frequencies.
@@ 3.3. Data Augmentation
Recently, a few studies have investigated data augmentation from a frequency domain perspective for time series (Wen0Y_augmentaion_2021, ). Since the frequency domain contains some vital information for time series analysis, such as periodic patterns, existing methods incorporate frequency domain features with time domain features for data augmentations with the aim of enhancing time series representations. For example, CoST (CoST_2022, ) incorporates a novel frequency domain contrastive loss which encourages discriminative seasonal representations and sidesteps the issue of determining the period of seasonal patterns present in the time series data. BTSF (Yang_ICML_2022, ) fuses the temporal and spectral features to enhance the discriminativity and expressiveness of the representations. TS-TFC (LiuMMW23, ) proposes a temporal-frequency co-training model for time-series semi-supervised learning, utilizing the complementary information from two distinct views for unlabeled data learning.
More recently, different from CoST and BTSF that apply DFT after augmenting samples in the time domain, one new approach named TF-C (zhang_nips_2022, ) introduces frequency domain augmentations that directly perturb the frequency spectrum. It develops frequency-based contrastive augmentation to leverage rich spectral information and directly perturbs the frequency spectrum to leverage frequency-invariance for contrastive learning. Compared to performing data augmentations directly in the frequency domain (e.g., TF-C), applying the FFT after augmenting samples in the time domain (e.g., CoST and BTSF) may lead to information loss.
@@ 3.4. Fourier Neural Operator Learning
According to the convolution theorem, differentiation is equivalent to multiplication in the Fourier domain (LiKALBSA2021, ). This efficiency property makes DFT frequently used to solve differential equations.
Recently, Fourier Neural Operators (FNOs) (LiKALBSA2021, ), which is currently the most promising one of the neural operators (Kovachki2021, ), have been proposed as an effective framework to solve partial differential equations (PDEs). More recently, FNO has been introduced in time series forecasting. FEDformer (fedformer_2022, ) proposes Fourier-enhanced blocks and Wavelet-enhanced blocks to capture important structures in time series through frequency domain mapping. FourierGNN (EVFGN_2022, ) reformulates the graph convolution operator in the frequency domain and efficiently computes graph convolutions over a hypervariate graph which represents the high-resolution correlations between any two variables at any two timestamps.
@@ 4.1. Complex-Value Data
The DFT output values are complex and can be represented in two ways. One representation is through the real and imaginary parts (as shown in Equation (2)), while the other representation is through the amplitude and phase parts (as depicted in Equations (3) and (4)). While it is possible to simplify the calculation by retaining only one part, such as discarding the imaginary components (GodfreyG_tnnls_2018, ), this approach may result in information loss.
In fact, there are two main approaches for performing neural networks on complex values. One approach is to treat each part of the complex value as a feature and then feed them to neural networks, respectively. Afterward, the output of corresponding networks is combined as a complex type (e.g., like Equation (2)), and then the inverse DFT is executed and transmitted to the time domain. For example, StemGNN (StemGNN_2020, ) conducts GLU (glu2017, ) on real and imaginary parts, respectively, which concatenates them as a complex value and applies IDFT. ATFN (ATFN_2022, ) utilizes two linear layers to process the amplitude and phase parts, respectively, and then combines them as a whole. The other is to conduct complex multiplication in the frequency domain directly. FEDformer (fedformer_2022, ) randomly samples a few frequencies and conducts complex multiplication with a parameterized kernel incorporated with attention architecture.
@@ 4.2. Real-Value Data
The output value type of DCT and DWT is real, hence commonly used network structures can be directly applied to them. For instance, DEPTS (fanwei2022, ) performs DCT and can be seamlessly integrated with an MLP to process the output parameters of DCT. Although the output value type of DFT is complex, some work discards one part, such as the phase part (ZhangAQ_kdd_2017, ), making their network design effectively real-valued. Additionally, some methods just apply filtering in the frequency domain and transform back to the time domain, resulting in network designs that also belong to real-valued networks.
Except for capturing frequency patterns, in contrast to other network designs, one main purpose of network design for frequency-based models is the frequency component selection to decide which component is discriminative or critical. For example, RobustPeriod (Wen_sigmod_2021, ) applies DWT to decouple time series into multiple levels of wavelet coefficients and then proposes a method to robustly calculate unbiased wavelet variance at each level and rank periodic possibilities. TimesNet (timesnet_2023, ) selects the top-$k$ amplitude values while discarding the remaining ones. FAN (fan_2024, ) decomposes frequency components into non-stationary and stationary parts and employs specialized network modules for each.
@@ 5. Applications
SFM (ZhangAQ_kdd_2017, ) DFT Feature engineering Real-value RNN Forecasting Decomposition
StemGNN (StemGNN_2020, ) DFT Feature engineering Complex-value GLU Forecasting Decomposition
Autoformer (autoformer_2021, ) DFT Feature engineering Complex-value Transformer Forecasting Global view Efficiency
DEPTS (fanwei2022, ) DCT Feature engineering Real-value MLP Forecasting Decomposition
FEDformer (fedformer_2022, ) DFT Feature engineering Operator learning Complex-value Transformer Forecasting Global view Efficiency
CoST (CoST_2022, ) DFT Data augmentation Complex-value MLP Forecasting Decomposition
FiLM (FiLM_2022, ) DFT Compression Complex-value MLP Forecasting Sparse Representation
FreDo (fredo2022, ) DFT Feature engineering Complex-value MLP Forecasting Decompostion
WAVEFORM (YangLWZPW23, ) DWT Feature engineering Real-value GCN Forecasting Decompostion
FourierGNN (EVFGN_2022, ) DFT Operator learning Complex-value MLP Forecasting Efficiency
FreTS (frets_2023, ) DFT Operator learning Complex-value MLP Forecasting Global view
FilterNet (filternet_2024, ) DFT Feature engineering Complex-value MLP Forecasting Decomposition
FAN (fan_2024, ) DFT Feature engineering Real-value MLP Forecasting Decomposition
SR-CNN (Ren_KDD_2019, ) DFT Feature engineering Real-value CNN Anomaly Detection Decomposition
RobustTAD (Gao_2020, ) DFT Data augmentation Complex-value CNN Anomaly Detection Decomposition
TFAD (TFAD_2022, ) DWT Feature engineering Real-value TCN Anomaly Detection Decomposition
FITS (fits_2024, ) DFT Compression Complex-value MLP Forecasting Anomaly Detection Sparse Representation
RCF (WangWLW_kdd_2018, ) DWT Feature engineering Real-value CNN Classification Decomposition
WD (Khan_nips_2018, ) DWT Feature engineering Real-value CNN Classification Decomposition
BTSF (Yang_ICML_2022, ) DFT Data augmentation Real-value CNN Classification Forecasting Decomposition
TF-C (zhang_nips_2022, ) DFT Data augmentation Real-value Transformer Classification Decomposition
TimesNet (timesnet_2023, ) DFT Feature engineering Real-value CNN Classification Forecasting Anomaly Detection Imputation Decompostion
FGTI (fgti_2024, ) DFT Feature engineering Complex-value Transformer Imputation Decompostion
PSW-I (wang2025optimal, ) DFT Feature engineering Complex-value MLP Imputation Decompostion
@@ 5.1. Time Series Forecasting
Time series forecasting is essential in various domains, such as decision making and financial analysis. Recently, some methods leverage frequency information to improve the accuracy or efficiency of time series forecasting. SFM (ZhangAQ_kdd_2017, ) decomposes the hidden states of memory cells into multiple frequency components and models multi-frequency trading patterns. StemGNN (StemGNN_2020, ) learns spectral representations which are easier to recognize after DFT. Autoformer (autoformer_2021, ) leverages FFT to calculate auto-correlation efficiently. DEPTS (fanwei2022, ) conducts DCT to extract periodic features and then applies multi-layer perceptrons to these features for periodicity dependencies in time series. FEDformer (fedformer_2022, ) captures the global view of time series in the frequency domain. CoST (CoST_2022, ) learns the seasonal representations in the frequency domain. FiLM (FiLM_2022, ) utilizes Fourier analysis to keep low-frequency Fourier components. DeRiTS (wei_derits_2024, ) and FAN (fan_2024, ) address the non-stationary issues for time series forecasting from the frequency spectrum perspective. FreDF (fredf_2025, ) mitigates label correlation by learning to forecast in the frequency domain, thereby reducing estimation bias.
@@ 5.2. Time Series Anomaly Detection
Recently, frequency-based models have been introduced in anomaly detection. TFAD (TFAD_2022, ) takes advantage of frequency domain analysis for seasonality anomaly. MACE (mace_2024, ) introduces a pattern extraction mechanism that exploits the inherent sparsity of the frequency domain to improve the model’s generalization across diverse normal patterns. CATCH (wu2025catch, ) patchifys the frequency domain into distinct frequency bands, and adaptively discover channel correlation across frequency bands to effectively detect both point and subsequence anomalies. FCVAE (FCVAE_2024, ) integrates global and local frequency features to capture long-periodic heterogeneous patterns and fine-grained short-periodic trends, enabling effective unsupervised anomaly detection in univariate time series. TSAD (TSAD_2024, ) employs nested sliding windows, where the outer and inner windows correspond to the time and frequency domains to bridge the discrepancy between time and frequency representations.
@@ 5.3. Time Series Classification
Time series classification is an important and challenging problem in time series analysis. Recently, a few models have considered frequency domain information to perform this task. RCF (WangWLW_kdd_2018, ) extracts distinguishing features from the DWT decomposed results. WD (Khan_nips_2018, ) uses wavelet functions with adjustable scale parameters to learn the spectral decomposition directly from the signal. BTSF (Yang_ICML_2022, ) fuses time and spectral information to enhance the discriminativity and expressiveness of the representations. TF-C (zhang_nips_2022, ) develops frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency in time series. TSLANet (TSLANet_2024, ) leverages the power of Fourier transform alongside global and local filters to cover the whole frequency spectrum, while adaptively removing high frequencies that tend to introduce noises.
@@ 5.4. Time Series Imputation
Time series imputation refers to the process of filling in missing values in a time-dependent dataset. FGTI (fgti_2024, ) employs a frequency-aware diffusion model that uses high-frequency filters for residual imputation and dominant-frequency filters for trend and seasonal refinement. PSW-I (wang2025optimal, ) This work presents a novel Proximal Spectrum Wasserstein (PSW) discrepancy, which integrates pairwise spectral distance and selective matching regularization to accurately quantify distributional discrepancies between two sets of time series, specifically tailored for the imputation task. Both methods incorporate frequency technology to enhance the accuracy of imputations by addressing different components of the time series.
@@ 6.2. Limitations
Frequency transformation techniques, including DFT and DCT, primarily emphasize capturing the frequency characteristics of a time series. While these techniques offer valuable insights into the frequency domain, they may overlook or inadequately represent temporal information (yang2024rethinking, ). Certain temporal patterns or dynamics inherent in the time series may not be fully captured in the frequency domain, thereby limiting the comprehensive analysis and understanding of the temporal aspects (GodfreyG_tnnls_2018, ).
@@
Frequency transformation techniques often require setting parameters, such as window size, sampling rate, or frequency bands. Selecting appropriate parameter values can be challenging, and suboptimal choices may lead to inaccurate frequency representations or missing important frequency components (Khan_nips_2018, ; pnas_2022, ). Accordingly, parameter tuning and optimization are necessary to ensure the effectiveness of frequency transformation in time series analysis.
@@ 7.1.1. Leveraging New Orthogonal Transform Technology
Recent studies have shown the efficiency and effectiveness of orthogonal transform which serves as a plug-in operation in neural networks, including frequency analysis and polynomial family. Some new orthogonal transform technologies have been introduced in neural networks and achieved good results. For example, the Fractional Fourier transform (FrFT) has been proven to be desirable for noise removal and can enhance the discrimination between anomalies and background (TaoZLLD_2019, ). In (ZhaoTLPL_2022, ), the authors utilize FrFT to enhance efficient feature fusion and comprehensive feature extraction. (zhao2022, ) leverages FrFT to enable flexible extraction of global contexts and sequential spectral information. In the future, it would be a promising direction to incorporate more new orthogonal transform technologies for deep learning in time series analysis, such as FrFT.

### v6

@@ 3.2. Compression
Previous works utilize frequency transformation to obtain sparse representations and remove redundant information in the frequency domain. Moreover, since noise signals usually appear as high frequencies, it is easy to filter them out in the frequency domain. For example, in FiLM (Zhou et al., 2022a), authors view time series forecasting from the sequence compression perspective and apply Fourier analysis to keep the part of the representation related to low-frequency Fourier components to remove the impact of noises. (Rippel et al., 2015) proposes spectral pooling that performs dimensionality reduction by truncating the representation in the frequency domain because energy is heavily concentrated in the lower frequencies. (Xu et al., 2020) proposes a learning-based frequency selection method to identify the trivial frequency components while removing redundant information.
@@ 3.3. Data Augmentation
Recently, a few studies have investigated data augmentation from a frequency domain perspective for time series (Wen et al., 2021b). Since the frequency domain contains some vital information for time series analysis, such as periodic patterns, existing methods incorporate frequency domain features with time domain features for data augmentations with the aim of enhancing time series representations. For example, CoST (Woo et al., 2022a) incorporates a novel frequency domain contrastive loss which encourages discriminative seasonal representations and sidesteps the issue of determining the period of seasonal patterns present in the time series data. BTSF (Yang and Hong, 2022) fuses the temporal and spectral features to enhance the discriminativity and expressiveness of the representations. TS-TFC (Liu et al., 2023) proposes a temporal-frequency co-training model for time-series semi-supervised learning, utilizing the complementary information from two distinct views for unlabeled data learning.
More recently, different from CoST and BTSF that apply DFT after augmenting samples in the time domain, one new approach named TF-C (Zhang et al., 2022a) introduces frequency domain augmentations that directly perturb the frequency spectrum. It develops frequency-based contrastive augmentation to leverage rich spectral information and directly perturbs the frequency spectrum to leverage frequency-invariance for contrastive learning. Compared to performing data augmentations directly in the frequency domain (e.g., TF-C), applying the FFT after augmenting samples in the time domain (e.g., CoST and BTSF) may lead to information loss.
@@ 3.4. Fourier Neural Operator Learning
According to the convolution theorem, differentiation is equivalent to multiplication in the Fourier domain (Li et al., 2021). This efficiency property makes DFT frequently used to solve differential equations.
Recently, Fourier Neural Operators (FNOs) (Li et al., 2021), which is currently the most promising one of the neural operators (Kovachki et al., 2021), have been proposed as an effective framework to solve partial differential equations (PDEs). More recently, FNO has been introduced in time series forecasting. FEDformer (Zhou et al., 2022b) proposes Fourier-enhanced blocks and Wavelet-enhanced blocks to capture important structures in time series through frequency domain mapping. FourierGNN (Yi et al., 2023a) reformulates the graph convolution operator in the frequency domain and efficiently computes graph convolutions over a hypervariate graph which represents the high-resolution correlations between any two variables at any two timestamps.
@@ 4.1. Complex-Value Data
The DFT output values are complex and can be represented in two ways. One representation is through the real and imaginary parts (as shown in Equation (2)), while the other representation is through the amplitude and phase parts (as depicted in Equations (3) and (4)). While it is possible to simplify the calculation by retaining only one part, such as discarding the imaginary components (Godfrey and Gashler, 2018), this approach may result in information loss.
In fact, there are two main approaches for performing neural networks on complex values. One approach is to treat each part of the complex value as a feature and then feed them to neural networks, respectively. Afterward, the output of corresponding networks is combined as a complex type (e.g., like Equation (2)), and then the inverse DFT is executed and transmitted to the time domain. For example, StemGNN (Cao et al., 2020) conducts GLU (Dauphin et al., 2017) on real and imaginary parts, respectively, which concatenates them as a complex value and applies IDFT. ATFN (Yang et al., 2022) utilizes two linear layers to process the amplitude and phase parts, respectively, and then combines them as a whole. The other is to conduct complex multiplication in the frequency domain directly. FEDformer (Zhou et al., 2022b) randomly samples a few frequencies and conducts complex multiplication with a parameterized kernel incorporated with attention architecture.
@@ 4.2. Real-Value Data
The output value type of DCT and DWT is real, hence commonly used network structures can be directly applied to them. For instance, DEPTS (Fan et al., 2022) performs DCT and can be seamlessly integrated with an MLP to process the output parameters of DCT. Although the output value type of DFT is complex, some work discards one part, such as the phase part (Zhang et al., 2017), making their network design effectively real-valued. Additionally, some methods just apply filtering in the frequency domain and transform back to the time domain, resulting in network designs that also belong to real-valued networks.
Except for capturing frequency patterns, in contrast to other network designs, one main purpose of network design for frequency-based models is the frequency component selection to decide which component is discriminative or critical. For example, RobustPeriod (Wen et al., 2021a) applies DWT to decouple time series into multiple levels of wavelet coefficients and then proposes a method to robustly calculate unbiased wavelet variance at each level and rank periodic possibilities. TimesNet (Wu et al., 2023) selects the top-$k$ amplitude values while discarding the remaining ones. FAN (Ye et al., 2024) decomposes frequency components into non-stationary and stationary parts and employs specialized network modules for each.
@@ 5. Applications
SFM (Zhang et al., 2017) DFT Feature engineering Real-value RNN Forecasting Decomposition
StemGNN (Cao et al., 2020) DFT Feature engineering Complex-value GLU Forecasting Decomposition
Autoformer (Wu et al., 2021) DFT Feature engineering Complex-value Transformer Forecasting Global view Efficiency
DEPTS (Fan et al., 2022) DCT Feature engineering Real-value MLP Forecasting Decomposition
FEDformer (Zhou et al., 2022b) DFT Feature engineering Operator learning Complex-value Transformer Forecasting Global view Efficiency
CoST (Woo et al., 2022a) DFT Data augmentation Complex-value MLP Forecasting Decomposition
FiLM (Zhou et al., 2022a) DFT Compression Complex-value MLP Forecasting Sparse Representation
WAVEFORM (Yang et al., 2023) DWT Feature engineering Real-value GCN Forecasting Decompostion
FourierGNN (Yi et al., 2023a) DFT Operator learning Complex-value MLP Forecasting Efficiency
FreTS (Yi et al., 2023b) DFT Operator learning Complex-value MLP Forecasting Global view
FilterNet (Yi et al., 2024a) DFT Feature engineering Complex-value MLP Forecasting Decomposition
FreDF (Wang et al., 2025a) DFT Feature engineering Complex-value MLP Forecasting Decomposition
SR-CNN (Ren et al., 2019) DFT Feature engineering Real-value CNN Anomaly Detection Decomposition
RobustTAD (Gao et al., 2020) DFT Data augmentation Complex-value CNN Anomaly Detection Decomposition
TFAD (Zhang et al., 2022b) DWT Feature engineering Real-value TCN Anomaly Detection Decomposition
FITS (Xu et al., 2024) DFT Compression Complex-value MLP Forecasting Anomaly Detection Sparse Representation
RCF (Wang et al., 2018) DWT Feature engineering Real-value CNN Classification Decomposition
WD (Khan and Yener, 2018) DWT Feature engineering Real-value CNN Classification Decomposition
BTSF (Yang and Hong, 2022) DFT Data augmentation Real-value CNN Classification Forecasting Decomposition
TF-C (Zhang et al., 2022a) DFT Data augmentation Real-value Transformer Classification Decomposition
TimesNet (Wu et al., 2023) DFT Feature engineering Real-value CNN Classification Forecasting Anomaly Detection Imputation Decompostion
FGTI (Yang et al., 2024b) DFT Feature engineering Complex-value Transformer Imputation Decompostion
PSW-I (Wang et al., 2025b) DFT Feature engineering Complex-value MLP Imputation Decompostion
@@ 5.1. Time Series Forecasting
Time series forecasting is essential in various domains, such as decision making and financial analysis. Recently, some methods leverage frequency information to improve the accuracy or efficiency of time series forecasting. SFM (Zhang et al., 2017) decomposes the hidden states of memory cells into multiple frequency components and models multi-frequency trading patterns. StemGNN (Cao et al., 2020) learns spectral representations which are easier to recognize after DFT. Autoformer (Wu et al., 2021) leverages FFT to calculate auto-correlation efficiently. DEPTS (Fan et al., 2022) conducts DCT to extract periodic features and then applies multi-layer perceptrons to these features for periodicity dependencies in time series. FEDformer (Zhou et al., 2022b) captures the global view of time series in the frequency domain. CoST (Woo et al., 2022a) learns the seasonal representations in the frequency domain. FiLM (Zhou et al., 2022a) utilizes Fourier analysis to keep low-frequency Fourier components. DeRiTS (Fan et al., 2024) and FAN (Ye et al., 2024) address the non-stationary issues for time series forecasting from the frequency spectrum perspective. FreDF (Wang et al., 2025a) mitigates label correlation by learning to forecast in the frequency domain, thereby reducing estimation bias.
@@ 5.2. Time Series Anomaly Detection
Recently, frequency-based models have been introduced in anomaly detection. RobustTAD (Gao et al., 2020) explores the data augmentation methods in the frequency domain to further increase labeled data. TFAD (Zhang et al., 2022b) takes advantage of frequency domain analysis for seasonality anomaly. MACE (Chen et al., 2024) introduces a pattern extraction mechanism that exploits the inherent sparsity of the frequency domain to improve the model’s generalization across diverse normal patterns. CATCH (Wu et al., 2025) patchifys the frequency domain into distinct frequency bands, and adaptively discover channel correlation across frequency bands to effectively detect both point and subsequence anomalies. FCVAE (Wang et al., 2024a) integrates global and local frequency features to capture long-periodic heterogeneous patterns and fine-grained short-periodic trends, enabling effective unsupervised anomaly detection in univariate time series. TSAD (Nam et al., 2024) employs nested sliding windows, where the outer and inner windows correspond to the time and frequency domains to bridge the discrepancy between time and frequency representations.
@@ 5.3. Time Series Classification
Time series classification is an important and challenging problem in time series analysis. Recently, a few models have considered frequency domain information to perform this task. RCF (Wang et al., 2018) extracts distinguishing features from the DWT decomposed results. WD (Khan and Yener, 2018) uses wavelet functions with adjustable scale parameters to learn the spectral decomposition directly from the signal. BTSF (Yang and Hong, 2022) fuses time and spectral information to enhance the discriminativity and expressiveness of the representations. TF-C (Zhang et al., 2022a) develops frequency-based contrastive augmentation to leverage rich spectral information and explore time-frequency consistency in time series. TSLANet (Eldele et al., 2024) leverages the power of Fourier transform alongside global and local filters to cover the whole frequency spectrum, while adaptively removing high frequencies that tend to introduce noises.
@@ 5.4. Time Series Imputation
Time series imputation refers to the process of filling in missing values in a time-dependent dataset. FGTI (Yang et al., 2024b) employs a frequency-aware diffusion model that uses high-frequency filters for residual imputation and dominant-frequency filters for trend and seasonal refinement. PSW-I (Wang et al., 2025b) This work presents a novel Proximal Spectrum Wasserstein (PSW) discrepancy, which integrates pairwise spectral distance and selective matching regularization to accurately quantify distributional discrepancies between two sets of time series, specifically tailored for the imputation task. Both methods incorporate frequency technology to enhance the accuracy of imputations by addressing different components of the time series.
@@ 6.2. Limitations
Frequency transformation techniques, including DFT and DCT, primarily emphasize capturing the frequency characteristics of a time series. While these techniques offer valuable insights into the frequency domain, they may overlook or inadequately represent temporal information (Yang et al., 2024a). Certain temporal patterns or dynamics inherent in time series may not be fully captured in the frequency domain, thereby limiting the comprehensive analysis and understanding of temporal aspects (Godfrey and Gashler, 2018).
@@
Frequency transformation techniques often require setting parameters, such as window size, sampling rate, or frequency bands. Selecting appropriate parameter values can be challenging, and suboptimal choices may lead to inaccurate frequency representations or missing important frequency components (Khan and Yener, 2018; Michau et al., 2022). Accordingly, parameter tuning and optimization are necessary to ensure the effectiveness of frequency transformation in time series analysis.
@@ 7.1.1. Leveraging New Orthogonal Transform Technology
Recent studies have shown the efficiency and effectiveness of orthogonal transform which serves as a plug-in operation in neural networks, including frequency analysis and polynomial family. Some new orthogonal transform technologies have been introduced in neural networks and achieved good results. For example, (Park et al., 2021, 2024) propose the Partial Fourier Transform (PFT), an efficient and accurate algorithm designed to compute only a subset of Fourier coefficients, rather than the full spectrum. The Fractional Fourier transform (FrFT) has been proven to be desirable for noise removal and can enhance the discrimination between anomalies and background (Tao et al., 2019). In (Zhao et al., 2022a), the authors utilize FrFT to enhance efficient feature fusion and comprehensive feature extraction. (Zhao et al., 2022b) leverages FrFT to enable flexible extraction of global contexts and sequential spectral information. In the future, it would be a promising direction to incorporate more new orthogonal transform technologies for deep learning in time series analysis, such as FrFT.